OpenStudy (anonymous):

If y^2-2xy=21, then dy/dx at the point (2,-3) is

OpenStudy (anonymous):

implicit differentiation 2yy'+0xy-2y-2xy'=0 2yy'-2xy'=2y yy'-xy'=y y'(y-x)=y y'=y/(y-x)

OpenStudy (anonymous):

plug in your values

OpenStudy (anonymous):

can u explain how u do implicit diff. ?

OpenStudy (anonymous):

im kinda bad at it

OpenStudy (anonymous):

use the chain rule whenever you take the derivative of y with respect to x: for example, dy/dx (y^3)=3y^2y'

OpenStudy (anonymous):

so when u differentiate, u put y' to show it was differentiated?

OpenStudy (anonymous):

dy/dx y=y'

OpenStudy (anonymous):

the chain rule is u multiply derivative of inside* outside + inside*derivative of outside, right?

OpenStudy (anonymous):

think you are talking about the product rule

OpenStudy (anonymous):

oh sorry so how do u use the chain rule?

OpenStudy (anonymous):

that wasn't a very good explanation lol look here http://en.wikipedia.org/wiki/Chain_rule

OpenStudy (anonymous):

so how did u know to use teh chain rule becuase here it isnt a compostion of funtions

OpenStudy (anonymous):

think \(y=f(x)\)

OpenStudy (anonymous):

\[y^2-2xy=21\] \[f^2(x)-2xf(x)=521\]

OpenStudy (anonymous):

\[2f(x)f'(x)-2f(x)-2xf'(x)=0\] more easily written as \[2yy'-2y-2xy'=0\]

OpenStudy (anonymous):

ok i see that

OpenStudy (anonymous):